show that the function f:R2 →R2 given f(x,y)=(x-y, x+y) is linear.


1
Expert's answer
2021-08-09T07:56:07-0400

First

"f(\alpha(x, y)) = f(\alpha x, \alpha y) = (\alpha x - \alpha y, \alpha x + \alpha y)"

"= (\alpha(x-y) , \alpha(x+y)) = \alpha((x-y), (x+y)) = \alpha f(x, y)"

Second

"f(x_1, y_1) + f(x_2, y_2) = (x_1-y_1, x_1+y_1) + (x_2-y_2, x_2+y_2) = (x_1-y_1+x_2-y_2, x_1+y_1+x_2+y_2)"

"= ((x_1+x_2)-(y_1+y_2), (x_1+x_2) + (y_1+y_2)) = f((x_1+x_2), (y_1+y_2))"

So f is a linear transformation